# Forget BIT, It is All about TOKEN: Towards Semantic Information Theory for LLMs

### v1

@@ V-B Transformer Architecture
Consider a decomposition of $\mathbf{A}_{tj}$ as follows:
$\mathbf{A}_{tj}=\pi_{tj}\mathbf{A},$ (45)
where $\mathbf{A}$ is a time-invariant parameter matrix, and $\pi_{tj}$ is the only time-variant scalar weight satisfying $\sum_{j=1}^{t-1}\pi_{tj}=1$ and $\pi_{tj}\geq 0$. Simple derivation yields the following theorem.
@@
$\mathbf{u}_{t}=\arg\mathrm{softmax}\left(\frac{1}{\Xi}\tilde{\mathbf{u}}_{1:N}^{T}\left(\sum_{j=1}^{t-1}\pi_{tj}\mathbf{A}\mathbf{u}_{j}\right)\right),\quad t=n+1,\ldots,T,$ (46)
where $\pi_{tj}$ is the output of the $\mathrm{softmax}$, that is
$\pi_{tj}=\frac{\exp(\mathbf{u}_{t-1}^{T}\mathbf{B}\mathbf{u}_{j})}{\sum_{i=1}^{t-1}\exp(\mathbf{u}_{t-1}^{T}\mathbf{B}\mathbf{u}_{i})},\quad j=1,\ldots,t-1.$ (47)
@@
Let $\mathbf{q}_{t}$, $\mathbf{k}_{t}$, and $\mathbf{v}_{t}$ be sample vectors of random variables $Q_{t}$, $K_{t}$, and $V_{t}$. The attention scheme in [28] implies
$\left\{\begin{aligned} &\mathbf{q}_{t}=\mathbf{W}_{q}\mathbf{u}_{t},\\ &\mathbf{k}_{t}=\mathbf{W}_{k}\mathbf{u}_{t},\\ &\mathbf{v}_{t}=\mathbf{W}_{v}\mathbf{u}_{t},\end{aligned}\right$. (48)
for $t=1,\ldots,T$. The output of the Transformer is
$\mathbf{u}_{t}=\arg\mathrm{softmax}\left(\frac{1}{\Xi}\tilde{\mathbf{u}}_{1:N}^{T}\left(\sum_{j=1}^{t-1}\pi_{tj}\mathbf{v}_{j}\right)\right),\quad t=n+1,\ldots,T,$ (49)
@@
$\pi_{tj}=\frac{\exp(\mathbf{q}_{t-1}^{T}\mathbf{k}_{j})}{\sum_{i=1}^{t-1}\exp(\mathbf{q}_{t-1}^{T}\mathbf{k}_{i})},\quad j=1,\ldots,t-1$ (50)
@@
This theorem shows that the Transformer is equivalent to a decomposition of $\mathbf{A}_{tj}$ as follows:
$\mathbf{A}_{tj}=\pi_{tj}\mathbf{A},$ (52)
where $\pi_{tj}$ measures the semantic relevance from $\mathbf{u}_{j}$ with $j=1,\ldots,t-1$ for predicting $\mathbf{u}_{t}$. In an utterance, the semantic relevance is asymmetric between different tokens. Recalling Section IV, the inner product is used to measure the correlations of token-level semantic. For the asymmetric semantic relevance in an utterance, the inner-product based bilinear form for predicting $\mathbf{u}_{t}$ is introduced as follows:
@@
where $\mathbf{B}\neq\mathbf{B}^{T}$ in general. $\pi_{tj}$ can then be assigned by using $\mathrm{softmax}$ as Eq. (47). According to Jaynes’ maximum entropy principle [89], the $\mathrm{softmax}$ is a probability assignment on discrete sample space that maximize the entropy with the constraint on the first order moment. Therefore, the obtained estimation of the semantic relevance is the one with the maximum uncertainty, i.e., the best achievable estimation in the worst case.
@@ V-C ELBO of the Transformer
The performance of AR-LLM can be analyzed from variational inference perspective. Similar to [90], $J$ is introduced as a latent variable defined on $\{1,\ldots,T\}$. $\pi_{tj}$ can then be seen as the probability that choosing the position $J=j$. Thus, the prediction of $U_{t}$ in Eq. (46) is the expectation over $J$ as follows:
$\mathbf{u}_{t}=\arg\mathrm{softmax}\left(\frac{1}{\Xi}\tilde{\mathbf{u}}_{1:N}^{T}\mathbb{E}_{J\sim Q(\cdot|U_{n+1:t-1},S_{1:n};\{\mathbf{A},\mathbf{B}\})}\{\mathbf{A}\mathbf{u}_{J}\}\right),\quad t=n+1,\ldots,T,$ (54)
@@
$Q(j|U_{n+1:t-1},S_{1:n};\{\mathbf{A},\mathbf{B}\})=\pi_{tj},\quad j=1,\ldots,t-1.$ (55)
@@
$\max_{\mathbf{A},\mathbf{B}}\mathrm{ELBO}(Q(J|U_{n+1:t-1}^{\hbar},S_{1:n};\{\mathbf{A},\mathbf{B}\})),\quad t=n+1,\ldots,T.$ (56)
@@
$\max_{\Phi}H(P_{t}^{\hbar},Q_{t}^{\Phi})=\min_{\Phi}\mathbb{E}_{P_{t}^{\hbar}}\{\log Q_{t}^{\Phi}\},\quad t=n+1,\ldots,T.$ (57)
@@
$Q_{t}^{\Phi^{\hbar}}=P(U_{t}|U_{n+1:t-1},S_{1:n};\Phi^{\hbar})=P(U_{t}^{\hbar}|U_{n+1:t-1}^{\hbar},S_{1:n})=P_{t}^{\hbar}.$ (58)
@@
$\log P(U_{t}^{h}|U_{n+1:t-1}^{\hbar},S_{1:n};\Phi)$ (60)
$=$ $\log\sum_{j=1}^{t-1}P(U_{t}^{h},j|U_{n+1:t-1}^{\hbar},S_{1:n};\Phi)$
$=$ $\log\sum_{j=1}^{t-1}P(U_{t}^{h},j|U_{n+1:t-1}^{\hbar},S_{1:n};\Phi)\frac{Q(j|U_{n+1:t-1}^{\hbar},S_{1:n};\{\mathbf{A},\mathbf{B}\})}{Q(j|U_{n+1:t-1}^{\hbar},S_{1:n};\{\mathbf{A},\mathbf{B}\})}$
$=$ $\log\mathbb{E}_{J\sim Q(\cdot|U_{n+1:t-1}^{\hbar},S_{1:n};\{\mathbf{A},\mathbf{B}\})}\left\{\frac{P(U_{t}^{\hbar},J|U_{n+1:t-1}^{\hbar},S_{1:n})}{Q(J|U_{n+1:t-1}^{\hbar},S_{1:n};\{\mathbf{A},\mathbf{B}\})}\right\}$
$\geq$ $\mathbb{E}_{J\sim Q(\cdot|U_{n+1:t-1}^{\hbar},S_{1:n};\{\mathbf{A},\mathbf{B}\})}\left\{\log\frac{P(U_{t}^{\hbar},J|U_{n+1:t-1}^{\hbar},S_{1:n})}{Q(J|U_{n+1:t-1}^{\hbar},S_{1:n};\{\mathbf{A},\mathbf{B}\})}\right\}.$
@@
$\mathrm{ELBO}(Q(J|U_{n+1:t-1}^{\hbar},S_{1:n};\{\mathbf{A},\mathbf{B}\}))$ (61)
$=$ $\mathbb{E}_{J\sim Q(\cdot|U_{n+1:t-1}^{\hbar},S_{1:n};\{\mathbf{A},\mathbf{B}\})}\{\log P(U_{t}^{\hbar},J|U_{n+1:T}^{\hbar},S_{1:n})\}$
$-D_{KL}(Q(J|U_{n+1:t-1}^{\hbar},S_{1:n};\{\mathbf{A},\mathbf{B}\})\|P(J|U_{n+1:t-1}^{\hbar},S_{1:n})).$
@@
$\max_{\mathbf{A},\mathbf{B}}\mathrm{ELBO}(Q(J|U_{n+1:t-1}^{\hbar},S_{1:n};\{\mathbf{A},\mathbf{B}\})),\quad t=n+1,\ldots,T.$ (62)
@@
$\max_{U_{t}\in\mathcal{S}(\Omega)}\mathrm{ELBO}(Q_{t}(J|U_{n+1:t-1},S_{1:n};\{\mathbf{A}^{\hbar+},\mathbf{B}^{\hbar+}\})),\quad t=n+1,\ldots,T,$ (63)
@@
In the inference phase, $U_{t}$ is chosen from $\mathcal{S}(\Omega)$ such that
$\log P(U_{t}|U_{n+1:t-1},S_{1:n};\Phi^{\hbar+})$ (64)
@@
$\log P(U_{t}|U_{n+1:t-1},S_{1:n};\Phi^{\hbar+})$ (65)
$=$ $\log\sum_{j=1}^{t-1}P(U_{t},j|U_{n+1:t-1},S_{1:n};\Phi^{\hbar+})$
$=$ $\log\sum_{j=1}^{t-1}P(U_{t},j|U_{n+1:t-1},S_{1:n};\Phi^{\hbar+})\frac{Q(j|U_{n+1:t-1},S_{1:n};\{\mathbf{A}^{\hbar+},\mathbf{B}^{\hbar+}\})}{Q(j|U_{n+1:t-1},S_{1:n};\{\mathbf{A}^{\hbar+},\mathbf{B}^{\hbar+}\})}$
$=$ $\log\mathbb{E}_{J\sim Q(\cdot|U_{n+1:t-1},S_{1:n};\{\mathbf{A}^{\hbar+},\mathbf{B}^{\hbar+}\})}\left\{\frac{P(U_{t},J|U_{n+1:t-1},S_{1:n};\Phi^{\hbar+})}{Q(J|U_{n+1:t-1},S_{1:n};\{\mathbf{A}^{\hbar+},\mathbf{B}^{\hbar+}\})}\right\}$
$\geq$ $\mathbb{E}_{J\sim Q(\cdot|U_{n+1:t-1},S_{1:n};\{\mathbf{A}^{\hbar+},\mathbf{B}^{\hbar+}\})}\left\{\log\frac{P(U_{t},J|U_{n+1:t-1},S_{1:n};\Phi^{\hbar+})}{Q(J|U_{n+1:t-1},S_{1:n};\{\mathbf{A}^{\hbar+},\mathbf{B}^{\hbar+}\})}\right\}.$
@@
$\mathrm{ELBO}(Q(J|U_{n+1:t-1},S_{1:n};\{\mathbf{A}^{\hbar+},\mathbf{B}^{\hbar+}\}))$ (66)
$=$ $\mathbb{E}_{J\sim Q(\cdot|U_{n+1:t-1},S_{1:n};\{\mathbf{A}^{\hbar+},\mathbf{B}^{\hbar+}\})}\{\log P(U_{t}|J,U_{n+1:t-1},S_{1:n};\Phi^{\hbar+})\}$
$-D_{KL}(Q(J|U_{n+1:t-1},S_{1:n};\{\mathbf{A}^{\hbar+},\mathbf{B}^{\hbar+}\})\|P(J|U_{n+1:t-1},S_{1:n};\Phi^{\hbar+})).$
@@
$\max_{U_{t}\in\mathcal{S}(\Omega)}\mathrm{ELBO}(Q(J|U_{n+1:t-1},S_{1:n};\{\mathbf{A}^{\hbar+},\mathbf{B}^{\hbar+}\})),\quad t=n+1,\ldots,T.$ (67)
@@ V-D Generalization Error Bound of the Transformer
Let $\mathbf{u}_{t}^{\hbar}$ be the ground-truth output vector at time $t$ for $t=n+1,\ldots,T$, where the corresponding random variable is $U_{t}^{\hbar}$. Therefore, the generalization error is given By
$H(P(U_{t}^{\hbar}),Q(U_{t})),$ (68)
where $P(U_{t}^{\hbar})$ is the one-shot coding, $Q(U_{t})$ is the output of the $\mathrm{softmax}$ function. Given $t$, we take $M$ samples from the Transformer output $U_{t}$, each of which is denoted as $\mathbf{u}_{mt}$ for $m=1,\ldots,M$. Recalling Theorem 5, the $i$-th entry of the logits $\mathbf{z}_{m}$ is defined by
$z_{m}^{i}=\frac{1}{\Xi}\tilde{\mathbf{u}}_{i}^{T}\left(\sum_{j=1}^{t-1}\pi_{tj}\mathbf{A}\mathbf{u}_{mj}\right),\quad i=1,\ldots,N.$ (69)
The empirical generalization error over a sample set with size $M$ is given by
$\hat{\mathcal{L}}(\mathbf{A},\mathbf{B})=\frac{1}{M}\sum_{m=1}^{M}\mathbf{1}^{T}(\mathbf{u}_{mt}^{\hbar})\log\frac{1}{\mathbf{q}(\mathbf{z}_{m})}=\frac{1}{M}\sum_{m=1}^{M}\log\frac{1}{q(z_{m}^{\hbar})},$ (70)
where $\mathbf{q}(\mathbf{z}_{m})$ is the output of the $\mathrm{softmax}$ function, and
$q(z_{m}^{\hbar})=\mathbf{1}^{T}(\mathbf{u}_{mt}^{\hbar})\mathbf{q}(\mathbf{z}_{m}).$ (71)
@@
$H(P(U_{t}^{\hbar}),Q(U_{t}))\leq\hat{\mathcal{L}}(\mathbf{A},\mathbf{B})+\frac{2\sqrt{2}}{M}\sum_{m=1}^{M}|z_{m}^{\hbar}|+3\sqrt{\frac{\log\frac{2}{\delta}}{2M}},\quad t=n+1,\ldots,T.$ (72)
with probability at least $1-\delta$ over the choice of $M$ samples.
@@
$\hat{\mathcal{R}}(\mathbf{A},\mathbf{B})=\mathbb{E}_{\bm{\sigma}}\left\{\sup_{\mathbf{A},\mathbf{B}}\frac{1}{M}\sum_{m=1}^{M}\sigma_{m}\log\frac{1}{q(z_{m}^{\hbar})}\right\},$ (73)
@@
$H(P(U_{t}^{\hbar}),Q(U_{t}))\leq\hat{\mathcal{L}}(\mathbf{A},\mathbf{B})+2\hat{\mathcal{R}}(\mathbf{A},\mathbf{B})+3\sqrt{\frac{\log\frac{2}{\delta}}{2M}}.$ (74)
Because $\mathbf{q}(\mathbf{z}_{m})$ is the output of the $\mathrm{softmax}$ function, $\hat{\mathcal{L}}(\mathbf{A},\mathbf{B})$ is $\sqrt{2}$-Lipschitz over $z_{m}^{\hbar}$ for $l^{2}$-norm. According to Talagrand’s Lemma in [52], we have
$\hat{\mathcal{R}}(\mathbf{A},\mathbf{B})\leq\mathbb{E}_{\bm{\sigma}}\left\{\sup_{\mathbf{A},\mathbf{B}}\frac{1}{M}\sum_{m=1}^{M}\sigma_{m}z_{m}^{\hbar}\right\}\leq\frac{\sqrt{2}}{M}\sum_{m=1}^{M}|z_{m}^{\hbar}|.$ (75)

### v2

@@ V-B Transformer Architecture
Consider a decomposition of $\mathbf{A}_{ij}$ as follows:
$\mathbf{A}_{ij}=\pi_{ij}\mathbf{A},$ (48)
where $\mathbf{A}$ is a time-invariant parameter matrix, and $\pi_{ij}$ is the only time-variant scalar weight satisfying $\sum_{j=1}^{i-1}\pi_{ij}=1$ and $\pi_{ij}\geq 0$. Simple derivation yields the following theorem.
@@
$\mathbf{u}_{i}=\arg\mathrm{softmax}\left(\frac{1}{T}\tilde{\mathbf{u}}_{1:M}^{T}\Psi_{\mathrm{FFN}}\left(\sum_{j=1}^{i-1}\pi_{ij}\mathbf{A}\mathbf{u}_{j}\right)\right),\quad i=n+1,\ldots,N,$ (49)
where $\Psi_{\mathrm{FFN}}(\cdot)$ is the feed-forward network, $\pi_{ij}$ is the output of the $\mathrm{softmax}$, that is
$\pi_{ij}=\frac{\exp(\mathbf{u}_{i-1}^{T}\mathbf{B}\mathbf{u}_{j})}{\sum_{l=1}^{i-1}\exp(\mathbf{u}_{i-1}^{T}\mathbf{B}\mathbf{u}_{l})},\quad j=1,\ldots,i-1.$ (50)
@@
Let $\mathbf{q}_{i}$, $\mathbf{k}_{i}$, and $\mathbf{v}_{i}$ be sample vectors of random variables $Q_{i}$, $K_{i}$, and $V_{i}$. The attention scheme in [28] implies
$\left\{\begin{aligned} &\mathbf{q}_{i}=\mathbf{W}_{q}\mathbf{u}_{i},\\ &\mathbf{k}_{i}=\mathbf{W}_{k}\mathbf{u}_{i},\\ &\mathbf{v}_{i}=\mathbf{W}_{v}\mathbf{u}_{i},\end{aligned}\right$. (51)
for $i=1,\ldots,N$. The output of the Transformer is
$\mathbf{u}_{i}=\arg\mathrm{softmax}\left(\frac{1}{T}\tilde{\mathbf{u}}_{1:M}^{T}\Psi_{\mathrm{FFN}}\left(\sum_{j=1}^{i-1}\pi_{ij}\mathbf{v}_{j}\right)\right),\quad i=n+1,\ldots,N,$ (52)
@@
$\pi_{ij}=\frac{\exp(\mathbf{q}_{i-1}^{T}\mathbf{k}_{j})}{\sum_{l=1}^{i-1}\exp(\mathbf{q}_{i-1}^{T}\mathbf{k}_{l})},\quad j=1,\ldots,i-1$ (53)
@@
This theorem shows that the Transformer is equivalent to a decomposition of $\mathbf{A}_{ij}$ as follows:
$\mathbf{A}_{ij}=\pi_{ij}\mathbf{A},$ (55)
where $\pi_{ij}$ measures the semantic relevance from $\mathbf{u}_{j}$ with $j=1,\ldots,i-1$ for predicting $\mathbf{u}_{i}$. In an utterance, the semantic relevance is asymmetric between different tokens. Recalling Section IV, the inner product is used to measure the correlations of token-level semantic. For the asymmetric semantic relevance in an utterance, the inner-product based bilinear form for predicting $\mathbf{u}_{i}$ is introduced as follows:
@@
where $\mathbf{B}\neq\mathbf{B}^{T}$ in general. $\pi_{ij}$ can then be assigned by using $\mathrm{softmax}$ as Eq. (50).
@@ V-C ELBO of the Transformer
The performance of AR-LLM can be analyzed from variational inference perspective. Similar to [94], $J$ is introduced as a latent variable defined on $\{1,\ldots,N\}$. $\pi_{ij}$ can then be seen as the probability that choosing the position $J=j$. Thus, the prediction of $U_{i}$ in Eq. (49) is the expectation over $J$ as follows:
$\mathbf{u}_{i}=\arg\mathrm{softmax}\left(\frac{1}{T}\tilde{\mathbf{u}}_{1:M}^{T}\Psi_{\mathrm{FFN}}\left(\mathbb{E}_{J\sim Q(\cdot|U_{n+1:i-1},S_{1:n};\{\mathbf{A},\mathbf{B}\})}\{\mathbf{A}\mathbf{u}_{J}\}\right)\right),\quad i=n+1,\ldots,N,$ (57)
@@
$Q(j|U_{n+1:i-1},S_{1:n};\{\mathbf{A},\mathbf{B}\})=\pi_{ij},\quad j=1,\ldots,i-1.$ (58)
@@
$\max_{\mathbf{A},\mathbf{B}}\mathrm{ELBO}(Q(J|U_{n+1:i-1}^{\hbar},S_{1:n};\{\mathbf{A},\mathbf{B}\})),\quad i=n+1,\ldots,N.$ (59)
@@
$\max_{\Phi}H(P_{i}^{\hbar},Q_{i}^{\Phi})=\min_{\Phi}\mathbb{E}_{P_{i}^{\hbar}}\{\log Q_{i}^{\Phi}\},\quad i=n+1,\ldots,N.$ (60)
@@
$Q_{i}^{\Phi^{\hbar}}=P(U_{i}|U_{n+1:i-1},S_{1:n};\Phi^{\hbar})=P(U_{i}^{\hbar}|U_{n+1:i-1}^{\hbar},S_{1:n})=P_{i}^{\hbar}.$ (61)
@@
$\log P(U_{i}^{\hbar}|U_{n+1:i-1}^{\hbar},S_{1:n};\Phi)$ (63)
$=$ $\log\sum_{j=1}^{i-1}P(U_{i}^{\hbar},j|U_{n+1:i-1}^{\hbar},S_{1:n};\Phi)$
$=$ $\log\sum_{j=1}^{i-1}P(U_{t}^{\hbar},j|U_{n+1:i-1}^{\hbar},S_{1:n};\Phi)\frac{Q(j|U_{n+1:i-1}^{\hbar},S_{1:n};\{\mathbf{A},\mathbf{B}\})}{Q(j|U_{n+1:i-1}^{\hbar},S_{1:n};\{\mathbf{A},\mathbf{B}\})}$
$=$ $\log\mathbb{E}_{J\sim Q(\cdot|U_{n+1:i-1}^{\hbar},S_{1:n};\{\mathbf{A},\mathbf{B}\})}\left\{\frac{P(U_{i}^{\hbar},J|U_{n+1:i-1}^{\hbar},S_{1:n})}{Q(J|U_{n+1:i-1}^{\hbar},S_{1:n};\{\mathbf{A},\mathbf{B}\})}\right\}$
$\geq$ $\mathbb{E}_{J\sim Q(\cdot|U_{n+1:i-1}^{\hbar},S_{1:n};\{\mathbf{A},\mathbf{B}\})}\left\{\log\frac{P(U_{i}^{\hbar},J|U_{n+1:i-1}^{\hbar},S_{1:n})}{Q(J|U_{n+1:i-1}^{\hbar},S_{1:n};\{\mathbf{A},\mathbf{B}\})}\right\}.$
@@
$\mathrm{ELBO}(Q(J|U_{n+1:i-1}^{\hbar},S_{1:n};\{\mathbf{A},\mathbf{B}\}))$ (64)
$=$ $\mathbb{E}_{J\sim Q(\cdot|U_{n+1:i-1}^{\hbar},S_{1:n};\{\mathbf{A},\mathbf{B}\})}\{\log P(U_{i}^{\hbar},J|U_{n+1:i-1}^{\hbar},S_{1:n})\}$
$-D_{KL}(Q(J|U_{n+1:i-1}^{\hbar},S_{1:n};\{\mathbf{A},\mathbf{B}\})\|P(J|U_{n+1:i-1}^{\hbar},S_{1:n})).$
@@
$\max_{\mathbf{A},\mathbf{B}}\mathrm{ELBO}(Q(J|U_{n+1:i-1}^{\hbar},S_{1:n};\{\mathbf{A},\mathbf{B}\})),\quad i=n+1,\ldots,N.$ (65)
@@
$\max_{U_{i}\in\mathcal{S}(\Omega)}\mathrm{ELBO}(Q_{i}(J|U_{n+1:i-1},S_{1:n};\{\mathbf{A}^{\hbar+},\mathbf{B}^{\hbar+}\})),\quad i=n+1,\ldots,N,$ (66)
@@
In the inference phase, $U_{i}$ is chosen from $\mathcal{S}(\Omega)$ such that
$\log P(U_{i}|U_{n+1:i-1},S_{1:n};\Phi^{\hbar+})$ (67)
@@
$\log P(U_{i}|U_{n+1:i-1},S_{1:n};\Phi^{\hbar+})$ (68)
$=$ $\log\sum_{j=1}^{i-1}P(U_{i},j|U_{n+1:i-1},S_{1:n};\Phi^{\hbar+})$
$=$ $\log\sum_{j=1}^{i-1}P(U_{i},j|U_{n+1:i-1},S_{1:n};\Phi^{\hbar+})\frac{Q(j|U_{n+1:i-1},S_{1:n};\{\mathbf{A}^{\hbar+},\mathbf{B}^{\hbar+}\})}{Q(j|U_{n+1:i-1},S_{1:n};\{\mathbf{A}^{\hbar+},\mathbf{B}^{\hbar+}\})}$
$=$ $\log\mathbb{E}_{J\sim Q(\cdot|U_{n+1:i-1},S_{1:n};\{\mathbf{A}^{\hbar+},\mathbf{B}^{\hbar+}\})}\left\{\frac{P(U_{i},J|U_{n+1:i-1},S_{1:n};\Phi^{\hbar+})}{Q(J|U_{n+1:i-1},S_{1:n};\{\mathbf{A}^{\hbar+},\mathbf{B}^{\hbar+}\})}\right\}$
$\geq$ $\mathbb{E}_{J\sim Q(\cdot|U_{n+1:i-1},S_{1:n};\{\mathbf{A}^{\hbar+},\mathbf{B}^{\hbar+}\})}\left\{\log\frac{P(U_{i},J|U_{n+1:i-1},S_{1:n};\Phi^{\hbar+})}{Q(J|U_{n+1:i-1},S_{1:n};\{\mathbf{A}^{\hbar+},\mathbf{B}^{\hbar+}\})}\right\}.$
@@
$\mathrm{ELBO}(Q(J|U_{n+1:i-1},S_{1:n};\{\mathbf{A}^{\hbar+},\mathbf{B}^{\hbar+}\}))$ (69)
$=$ $\mathbb{E}_{J\sim Q(\cdot|U_{n+1:i-1},S_{1:n};\{\mathbf{A}^{\hbar+},\mathbf{B}^{\hbar+}\})}\{\log P(U_{i}|J,U_{n+1:i-1},S_{1:n};\Phi^{\hbar+})\}$
$-D_{KL}(Q(J|U_{n+1:i-1},S_{1:n};\{\mathbf{A}^{\hbar+},\mathbf{B}^{\hbar+}\})\|P(J|U_{n+1:i-1},S_{1:n};\Phi^{\hbar+})).$
@@
$\max_{U_{i}\in\mathcal{S}(\Omega)}\mathrm{ELBO}(Q(J|U_{n+1:i-1},S_{1:n};\{\mathbf{A}^{\hbar+},\mathbf{B}^{\hbar+}\})),\quad i=n+1,\ldots,N.$ (70)
@@ V-D Generalization Error Bound of the Transformer
Let $\mathbf{u}_{i}^{\hbar}$ be the ground-truth output vector at time $i$ for $i=n+1,\ldots,N$, where the corresponding random variable is $U_{i}^{\hbar}$. Therefore, the generalization error is given By
$H(P(U_{i}^{\hbar}),Q(U_{i})),$ (71)
where $P(U_{i}^{\hbar})$ is the one-shot coding, $Q(U_{i})$ is the output of the $\mathrm{softmax}$ function. Given $i$, we take $K$ samples from the Transformer output $U_{i}$, each of which is denoted as $\mathbf{u}_{ki}$ for $k=1,\ldots,K$. Recalling Theorem 5, the $l$-th entry of the logits $\mathbf{z}_{k}$ is defined by
$z_{k}^{l}=\frac{1}{T}\tilde{\mathbf{u}}_{i}^{T}\Psi_{\mathrm{FFN}}\left(\sum_{j=1}^{i-1}\pi_{ij}\mathbf{A}\mathbf{u}_{kj}\right),\quad l=1,\ldots,M.$ (72)
The empirical generalization error over a sample set with size $K$ is given by
$\hat{\mathcal{L}}(\mathbf{A},\mathbf{B})=\frac{1}{K}\sum_{k=1}^{K}\mathbf{1}^{T}(\mathbf{u}_{ki}^{\hbar})\log\frac{1}{\mathbf{q}(\mathbf{z}_{k})}=\frac{1}{K}\sum_{k=1}^{K}\log\frac{1}{q(z_{k}^{\hbar})},$ (73)
where $\mathbf{q}(\mathbf{z}_{k})$ is the output of the $\mathrm{softmax}$ function, and
$q(z_{k}^{\hbar})=\mathbf{1}^{T}(\mathbf{u}_{ki}^{\hbar})\mathbf{q}(\mathbf{z}_{k}).$ (74)
@@
$H(P(U_{i}^{\hbar}),Q(U_{i}))\leq\hat{\mathcal{L}}(\mathbf{A},\mathbf{B})+\frac{2\sqrt{2}}{K}\sum_{k=1}^{K}|z_{k}^{\hbar}|+3\sqrt{\frac{\log\frac{2}{\delta}}{2K}},\quad i=n+1,\ldots,N.$ (75)
with probability at least $1-\delta$ over the choice of $K$ samples.
@@
$\hat{\mathcal{R}}(\mathbf{A},\mathbf{B})=\mathbb{E}_{\bm{\sigma}}\left\{\sup_{\mathbf{A},\mathbf{B}}\frac{1}{K}\sum_{k=1}^{K}\sigma_{k}\log\frac{1}{q(z_{k}^{\hbar})}\right\},$ (76)
@@
$H(P(U_{i}^{\hbar}),Q(U_{i}))\leq\hat{\mathcal{L}}(\mathbf{A},\mathbf{B})+2\hat{\mathcal{R}}(\mathbf{A},\mathbf{B})+3\sqrt{\frac{\log\frac{2}{\delta}}{2K}}.$ (77)
Because $\mathbf{q}(\mathbf{z}_{k})$ is the output of the $\mathrm{softmax}$ function, $\hat{\mathcal{L}}(\mathbf{A},\mathbf{B})$ is $\sqrt{2}$-Lipschitz over $z_{k}^{\hbar}$ for $l^{2}$-norm. According to Talagrand’s Lemma in [52], we have
$\hat{\mathcal{R}}(\mathbf{A},\mathbf{B})\leq\mathbb{E}_{\bm{\sigma}}\left\{\sup_{\mathbf{A},\mathbf{B}}\frac{1}{K}\sum_{k=1}^{K}\sigma_{k}z_{k}^{\hbar}\right\}\leq\frac{\sqrt{2}}{K}\sum_{k=1}^{K}|z_{k}^{\hbar}|.$ (78)